Thanks for helping me sole the equations. I sent an equation with error.
this is the correct equations.
Solve for A and B in the following problem.
4A+B=16, 2A+3B=24
Thank you!

Answer by gsmani_iyer(201)


4a + b = 16 ... (1)

2a + 3b = 24 ... (2)

Eq.(1)*3, we get

12a + 3b = 48 ... (3)
Eq.(3)-(2)

10a = 24
so, a = 2.4. By substituting this value of a in eq.no.1, we get

4*2.4 + b = 16

b = 16 - (4*2.4) i.e = 16 - 9.6

b = 6.4
a = 2.4; b = 6.4; Answer.
Is it clear now?
gsm